Wednesday, March 18, 2020, 2:00 pm EDT
From Hall A Wiki
*********BLUEJEANS CALL-IN INSTRUCTIONS********* ---------------------------------------------------- 1. Dial Toll-Free Number: 1-888-240-2560 (U.S. & Canada) 2. International participants dial: Toll Number: 1-408-740-7256 3. Enter meeting ID:140287970 Desktop sharing: https://bluejeans.com/140287970
Meeting room = None. Remote only.
- Keep taking data!
- QWAD preparation progress
- Analysis updates
- Laser polarization systematic tests: Preliminary results from HWP=73 and 56 deg. , 
- Other options for "reasonable" values of RRPD: 
- Compton edge decreasing 
- Steering vs. spectrum 
- Spectrum fitting for collimator offset: Target dependence? Drifting gain + low amplitude backgrounds?
- LED run
- Online analysis results
- Polarization vs. time
Mark D, Dave G, Adam Z, Kent P, Brian Q, JCC, Cip G, Amali P
- Photon Detector
- Table scan: Epics logger stopped writing to the CODA file so the scan is very difficult to recover. Should be redone after procedures are made more robust.
- Planned systematic study of angles of the electron beam will be done ASAP after the Moller today.
- Possible new additional background from calcium that was not observed earlier. Amali investigating.
- Tube gain is decreasing but still completely functional. Increasing the voltage to recover the gain might affect the linearity.
- Pulser trigger doesn't work when there is beam. Need someone to verify the trigger logic in the hall. Opportunistic access.
- Systematic checks wen't well. Other systematic tests are the "second solution" of the same sign or to change the sign. elog 575
- Electron Detector
- QWADs in EEL126 for channel adjustment.
- Possible access to reinstate old system.